a super-bouncy ball rebounds half the height it drops. the ball is dropped from a height of 176 feet. How far off the ground is the ball when it has travled a total of 500 feet?

Answer by rothauserc(4718)

The falls form a geometric series with initial term of 176 and a common ratio of (1/2)
:
The general form for the sum of a geometric series is
:
S(n) = a1 * (1-r^n) / (1-r)
:
For the falls, a(1) = 176 and r = (1/2)
:
S(n) = 176 * (1 - (1/2)^n) / (1 - (1/2))
:
For the rebounds, a(1) = 88 and r = (1/2)
:
S(n) = 88 * (1 - (1/2)^n) / (1 - (1/2))
:
We know
:
500 = (352 * (1 - (1/2)^(n+1))) + (176 * (1 - (1/2)^n))
:
divide both sides of = by 4
:
125 = (88 * (1 - (1/2)^(n+1))) + (44 * (1 - (1/2)^n))
:
125 = (88 * (1 - 2^(-n - 1))) + (44 * (1 - 2^(-n)))
:
7 * 2^n = 132
:
use definition of logarithm
:
n = 4.237
:
We calculate the rebounds for n = 4 and n = 3
:
S(4) = 88 * (1 - (1/2)^4) / (1 - (1/2)) = 165
S(3) = 88 * (1 - (1/2)^(4.237)) / (1 - (1/2)) = 154
:
***********************************************************
The ball is 165 - 154 = 11 feet above the ground
***********************************************************
: